OpenStudy (anonymous):

The line that passes through (1, 1) and (3, 4)

OpenStudy (anonymous):

so in other words you're looking for an equation of the line that passes through (1, 1) and (3, 4)??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first find the slope using y2-y1/x2-x1

OpenStudy (anonymous):

then use tangent line formula of: y-y1 = m(x-x1) plug the points. m is the slope

OpenStudy (anonymous):

y=3/2x-1

OpenStudy (anonymous):

don't just give answer, try to leading them @nb209